1.An executive invests $25,000, some at 6% and some at 5% annual interest. If he receives an annual return of $1380, how much is invested at each rate?
2.A restaurant has fixed costs of $148.75 per day and an average unit cost of $4.75 for each meal served. If a typical meal costs $6, how many customers must eat at the restaurant each day for the owner to make a profit?

3.A machine to manufacture fasteners has a setup cost of $1,100 and a unit cost of $0.003 for each fastener manufactured. A newer machine has a setup cost of $1,450 but a unit cost of only $0.0025 for each fastener manufactured. Find the break point. (Round your answer to the nearest whole unit.)

Answer by mananth(16949)

1.An executive invests $25,000, some at 6% and some at 5% annual interest. If he receives an annual return of $1380, how much is invested at each rate?
Investment Part I 5.00% per annum ----x
Investment part II 6.00% per annum ----y
The sum of the investments is $25,000.00
The sum of individual interests = $1,380.00
x+y 25000 ------------------------1
5.00% x+6.00% y = $1,380.00
Multiply by 100
5x+6y = $138,000.00 --------2

Multiply (1) by -5
we get

-5x-5y =-125000.00

Add this to (2)

y = $13,000.00

divide by 1

y = $13,000.00 investment at 6.00%
Balance $12,000.00 investment at 5.00%
CHECK
$12,000.00 @ 5.00% $600.00
$13,000.00 @ 6.00% $780.00
Total ---------------- $1,380.00


2.A restaurant has fixed costs of $148.75 per day and an average unit cost of $4.75 for each meal served. If a typical meal costs $6, how many customers must eat at the restaurant each day for the owner to make a profit?

Let number of meals served per day be x
cost = 148.75 + 4.75x
Price of meal = 6
Amount collected = 6x
for break even 148.75+4.75x=6x
148.75=6x-4.75x
148.75=1.25x
148.75/1.25=x
x= 119
more than 119 meals have to be served to make a profit

3.A machine to manufacture fasteners has a setup cost of $1,100 and a unit cost of $0.003 for each fastener manufactured. A newer machine has a setup cost of $1,450 but a unit cost of only $0.0025 for each fastener manufactured. Find the break point. (Round your answer to the nearest whole unit.)
Old Machine set up cost = 1100
variable cost = 0.003
let number of fasteners produced be x
Cost = 1100+0.003x
New Machine
cost = 1450+0.0025x
For finding the breakeven between the machines
1100+0.003x=1450+0.0025x
0.003x-0.0025x=1450-1100
0.0005x=350
x= 350/0.0005
x= 700,000 fasteners
m.ananth@hotmail.ca